OpenStudy (anonymous):

Solve 2^(x + 3) = 7 I just submitted this and I got 0.94 and it was wrong :/ What did I do wrong?

OpenStudy (anonymous):

@lgbasallote HELP!

OpenStudy (jiteshmeghwal9):

log_{2}{7}=x+3

OpenStudy (anonymous):

\[b^x=A\iff x=\frac{\ln(A)}{\ln(b)}\]

OpenStudy (lgbasallote):

arent we over this?

OpenStudy (lgbasallote):

i told you 0.94 was wrong...

OpenStudy (anonymous):

I trusted my instincts! And it was wronggg!

OpenStudy (lgbasallote):

\[\large 2^{x+3} = 7\] change to log form \[\implies \log_2 7 = x+ 3\] subtract 3 from both sides \[(\log_2 7) - 3 = x\]

OpenStudy (lgbasallote):

does that help?

OpenStudy (anonymous):

Yes let me solve that

OpenStudy (anonymous):

I got -0.19?

OpenStudy (anonymous):

this is what you need \[x+3=\frac{\ln(7)}{\ln(2)}\] \[x=\frac{\ln(7)}{\ln(2)}-3\] whatever that is

OpenStudy (anonymous):

Let me put it in the calculator

OpenStudy (anonymous):

2.31242306964

OpenStudy (lgbasallote):

lol how'd that happen..

OpenStudy (lgbasallote):

you should get the same result from a while ago

OpenStudy (lgbasallote):

that \[\log_2 7 - 3 = x\]thing

OpenStudy (anonymous):

Oh oops

OpenStudy (lgbasallote):

so what did you get for \[\log_2 7 - 3\]

OpenStudy (anonymous):

-0.19

OpenStudy (lgbasallote):

yup

OpenStudy (anonymous):

Ohhh for some reason I put in another answer even though that was was right. Nooo idea why.

OpenStudy (anonymous):

?

OpenStudy (anonymous):

^^^^ yes?

OpenStudy (anonymous):

Did you get different satellite

OpenStudy (lgbasallote):

if he did..he's probably wrong =))) hahaha

OpenStudy (anonymous):

no it is right my real question is, why write \(x=\log_2(7)-3\) ? i mean that is certainly correct, but unless you have an extra fancy calculator with log base two on it, you need to compute \[\frac{\ln(7)}{\ln(2)}-3\]

OpenStudy (anonymous):

^^^ Lol. Satellite does have a point! Much easier to calculate.

OpenStudy (lgbasallote):

well because ln's are usually used in calculus

OpenStudy (lgbasallote):

and it's too much latex work to do \[\frac{\log 7}{\log 2}\] rather than \[\log_2 7\]

OpenStudy (lgbasallote):

hayley knows what \(\log_2 7\) means anyway so heh..im using it..